A 30g sample of ore contains 25g of haematite, Fe2O3. What mass of iron could be extracted?
1
Expert's answer
2020-05-03T15:12:03-0400

m(Fe)=m(Fe2O3)*(2*Mr(Fe)/Mr(Fe2O3))=25*2*56/160=17.5 g

Answer: 17.5 g of iron could be extracted